# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ D.2 Useful Lemmas for Theorem 2
$\left|x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}\right|\leq\|x_{ti}\|_{H_{t}^{-1}}\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{2}}\leq\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}},$
where the first inequality is by Hölder’s inequality, and the last inequality holds by Lemma 1. Hence, it follows that
@@
D.2.2 Proof of Lemma D.2
Proof of Lemma D.2.
@@
To establish the proof for the second inequality, we return to (D.10):
@@
$\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right),$
@@
D.2.3 Proof of Lemma D.3
Proof of Lemma D.3.
@@
$\frac{\partial Q}{\partial i}=\frac{e^{u_{i}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
@@
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)+e^{u_{i}}e^{u_{j}}}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
$+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}\left(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}\right)}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{4}}$
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)\!+\!e^{u_{i}}e^{u_{j}}}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}}{(\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}})^{3}}.$
Let $p_{i}(\mathbf{u})=\frac{e^{u_{i}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ and $p_{0}(\mathbf{u})=\frac{\bar{v}_{0}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ For $i=j$, we have
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg{|}p_{i}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}-p_{i}(\mathbf{u})p_{j}(\mathbf{u})$
$+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
$=\bigg{|}p_{i}(\mathbf{u})p_{0}(\mathbf{u})-2p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
@@
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg{|}-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{\bar{v}_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
@@
Appendix E Proof of Lemma 1
@@
E.1 Main Proof of Lemma 1
@@
where ~$\tilde{\ell}_{t}(\mathbf{w})=\ell_{t}(\mathbf{w}_{t})+\langle\mathbf{w}-\mathbf{w}_{t},\nabla\ell_{t}(\mathbf{w}_{t})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{t}\|_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}$ and $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Let $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda>0$. Then, we have
@@
$-\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}.$ (E.1)
We first bound the first term in (E.1). For simplicity, we define the softmax function at round $t$ $\bm{\sigma}_{t}(\mathbf{z}):\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$ as follows:
$[\bm{\sigma}_{t}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})},\quad\forall i\in[K_{t}],$ (E.2)
where $[\cdot]_{i}$ denotes $i$’th element of the the input vector. We denote the probability of choosing the outside option as $[\bm{\sigma}_{t}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})}$. Although $[\bm{\sigma}_{t}(\mathbf{z})]_{0}$ is not the output of the softmax function, we represent it in a form similar to that in (E.2) for simplicity. Then, the user choice model in (1) can be expressed as $p_{t}(i|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{i}$ for all $i\in[K_{t}]$ and $p_{t}(0|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{0}$. Furthermore, the loss function (2) can also be written as $\ell(\mathbf{z}_{t},\mathbf{y}_{t})=\sum_{k=0}^{K_{t}}\mathbf{1}\left\{y_{ti}=1\right\}\cdot\log\left(\frac{1}{[\bm{\sigma}_{t}(\mathbf{z}_{t})]_{k}}\right)$.
@@
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})=\underbrace{\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})}_{(a)}+\underbrace{\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})}_{(b)},$ (E.3)
@@
Now, we are ready to prove the Lemma 1. By combining Lemma E.1, Lemma E.2, and Lemma E.3, we derive that
@@
where the second inequality holds because, by setting since $c=7\eta/6$ and $\lambda\geq\max\{84\sqrt{2}\eta,84d\eta\}$, we obtain:
@@
which conclude the proof of Lemma 1.
E.2 Proofs of Lemmas for Lemma 1
E.2.1 Proof of Lemma E.1
Proof of Lemma E.1.
Let ~$\tilde{\ell}_{s}(\mathbf{w})=\ell_{s}(\mathbf{w}_{s})+\langle\mathbf{w}-\mathbf{w}_{s},\nabla\ell_{s}(\mathbf{w}_{s})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{s}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s})}$ be a second-order approximation of the original function $\ell_{s}(\mathbf{w})$ at the point $\mathbf{w}_{s}$. The update rule (3) can also be expressed as
@@ E.2 Proofs of Lemmas for Lemma 1
Then, by Lemma E.4, we have
~$\langle\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle\leq\frac{1}{2\eta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$ (E.4)
To utilize Lemma E.6, we can rewrite the loss function as $\ell\left((x_{si}^{\top}\mathbf{w})_{i\in S_{s}},\mathbf{y}_{s}\right)=\ell_{s}(\mathbf{w})$. Consequently, according to Lemma E.6, it follows that
$\ell_{s}(\mathbf{w}_{s+1})-\ell_{s}(\mathbf{w}^{\star})\leq\langle\nabla\ell_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle-\frac{1}{\zeta}\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})},$ (E.5)
where $\zeta=\log(K+1)+4$. Then, by combining (E.4) and (E.5), we have
@@
where in the second equality, we apply the Taylor expansion by introducing $\bm{\xi}_{s+1}$, a convex combination of $\mathbf{w}_{s+1}$ and $\mathbf{w}_{s}$. The first inequality follows from Lemma B.1 and Proposition B.1, the second inequality holds by Assumption 1, and the last inequality holds because
@@
where the third equality is by setting $x_{s0}=\mathbf{0}$ for all $s\geq 1$.
@@
$\leq\zeta\left(\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})\right)+4\lambda+6\sqrt{2}\zeta\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2},$
where the first inequality holds since $\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{H_{1}}^{2}\leq\lambda\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{2}^{2}\leq 4\lambda$. Plugging in $\zeta=2\eta$, we conclude the proof. ∎
E.2.2 Proof of Lemma E.2
Proof of Lemma E.2.
@@
Note that $\tilde{\mathbf{z}}_{s}^{\mu}=\bm{\sigma}_{s}^{+}(\operatorname{smooth}^{\mu}_{s}(\bm{\sigma}_{s}(\tilde{\mathbf{z}}_{s})))$ by the definition of the pseudo inverse function $\bm{\sigma}_{s}^{+}$ such that $\bm{\sigma}_{s}^{+}(\bm{\sigma}_{s}(\mathbf{q}))=\mathbf{q}$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{s}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, by Lemma E.7, we have
$\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq 2\mu t,\quad\text{and}\quad\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu).$ (E.6)
@@
$=\sum_{s=1}^{t}\langle\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s},\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\rangle-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2},$ (E.7)
where $c_{\mu}=\log(K+1)+2\log(1+(K+1)/\mu)+2$, the inequality hols by Lemma E.6, and the last equality holds by a direct calculation of the first order and Hessian of the logistic loss as follows:
@@
Then, one can easily verify that $\|\mathbf{d}^{\prime}_{s}\|_{\infty}\leq 1$ since $\|\mathbf{z}_{s}^{\star}\|_{\infty}\leq\max_{i\in S_{s}}\|x_{xi}\|_{2}\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$. On the other hand, $\mathbf{d}_{s}^{\prime}$ is $\mathcal{F}_{s}$-measurable since $\mathbf{z}_{s}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$ are independent of $\mathbf{y}_{s}$. Moreover, we have $\|\mathbf{d}^{\prime}_{s}\|^{2}_{\mathbb{E}[\bm{\varepsilon}_{s}\bm{\varepsilon}_{s}^{\top}\mid\mathcal{F}_{s}]}=\|\mathbf{d}_{s}\|^{2}_{\mathbb{E}\left[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top}\mid\mathcal{F}_{s}\right]}=\|\mathbf{d}_{s}\|^{2}_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}$ and $\|\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}\|_{1}\leq 2$. Thus, by Lemma E.5, with probability at least $1-\delta$, for any $t\geq 1$, we have
@@
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right),$ (E.8)
where the second inequality holds because $\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}=\mathbf{d}_{s}^{\top}\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})\mathbf{d}_{s}\leq 2$ and $\lambda\geq 1$. Then, combining (E.8) and (E.7), and rearranging the terms, we obtain

### v6

@@ E.2 Proofs of Lemmas for Theorem 2
$\left|x_{ti}^{\top}\mathbf{w}_{t}-x_{ti}^{\top}\mathbf{w}^{\star}\right|\leq\|x_{ti}\|_{H_{t}^{-1}}\|\mathbf{w}_{t}-\mathbf{w}^{\star}\|_{H_{t}}\leq\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}},$
where the first inequality is by the Hölder’s inequality, and the last inequality holds by Lemma 1. Hence, it follows that
@@
E.2.2 Proof of Lemma E.2
Proof of Lemma E.2.
@@
To establish the proof for the second inequality, we return to (E.13):
@@
$\leq\frac{2}{\kappa}\sum_{s=1}^{t}\log\left(\frac{\det(H_{s+1})}{\det(H_{s})}\right)\leq\frac{2}{\kappa}d\log\left(1+\frac{t}{d\lambda}\right),$
@@
E.2.3 Proof of Lemma E.3
Proof of Lemma E.3.
@@
$\frac{\partial Q}{\partial i}=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
@@
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)+e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}$
$+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}\left(v_{0}+\sum_{k=1}^{K}e^{u_{k}}\right)}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{4}}$
$=\frac{\mathbbm{1}_{i=j}e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-\frac{e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}-\frac{\mathbbm{1}_{i=j}e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)\!+\!e^{u_{i}}e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{2}}+\frac{e^{u_{i}}\left(\sum_{k=1}^{K}e^{u_{k}}\right)2e^{u_{j}}}{(v_{0}+\sum_{k=1}^{K}e^{u_{k}})^{3}}.$
Let $p_{i}(\mathbf{u})=\frac{e^{u_{i}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ and $p_{0}(\mathbf{u})=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}$ For $i=j$, we have
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg{|}p_{i}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}-p_{i}(\mathbf{u})p_{j}(\mathbf{u})$
$+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
$=\bigg{|}p_{i}(\mathbf{u})p_{0}(\mathbf{u})-2p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
@@
$\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ $=\bigg{|}-p_{i}(\mathbf{u})p_{j}(\mathbf{u})-p_{i}(\mathbf{u})p_{j}(\mathbf{u})+2p_{i}(\mathbf{u})p_{j}(\mathbf{u})\frac{\sum_{k=1}^{K}e^{u_{k}}}{v_{0}+\sum_{k=1}^{K}e^{u_{k}}}\bigg{|}$
@@
Appendix F Proof of Lemma 1
@@
F.1 Main Proof of Lemma 1
@@
where ~$\tilde{\ell}_{t}(\mathbf{w})=\ell_{t}(\mathbf{w}_{t})+\langle\mathbf{w}-\mathbf{w}_{t},\nabla\ell_{t}(\mathbf{w}_{t})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{t}\|^{2}_{\nabla^{2}\ell_{t}(\mathbf{w}_{t})}$ and $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Let $\eta=\frac{1}{2}\log(K+1)+2$ and $\lambda>0$. Then, we have
@@
$-\sum_{i=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2}.$ (F.1)
We first bound the first term in (F.1). For simplicity, we define the softmax function at round $t$ $\bm{\sigma}_{t}(\mathbf{z}):\mathbb{R}^{K_{t}}\rightarrow\mathbb{R}^{K_{t}}$ as follows:
$[\bm{\sigma}_{t}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})},\quad\forall i\in[K_{t}],$ (F.2)
where $[\cdot]_{i}$ denotes $i$’th element of the the input vector. We denote the probability of choosing the outside option as $[\bm{\sigma}_{t}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K_{t}}\exp([\mathbf{z}]_{k})}$. Although $[\bm{\sigma}_{t}(\mathbf{z})]_{0}$ is not the output of the softmax function $\bm{\sigma}_{t}(\mathbf{z})$, we represent it in a form similar to that in (F.2) for simplicity. Then, the user choice model in (1) can be equivalently expressed as $p_{t}(i|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{i}$ for all $i\in[K_{t}]$ and $p_{t}(0|S_{t},\mathbf{w})=\left[\bm{\sigma}_{t}\left((x_{tj}^{\top}\mathbf{w})_{j\in S_{t}}\right)\right]_{0}$. Furthermore, the loss function (2) can also be written as $\ell(\mathbf{z}_{t},\mathbf{y}_{t})=\sum_{k=0}^{K_{t}}\mathbf{1}\left\{y_{ti}=1\right\}\cdot\log\left(\frac{1}{[\bm{\sigma}_{t}(\mathbf{z}_{t})]_{k}}\right)$.
@@
$\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})=\underbrace{\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})}_{(a)}+\underbrace{\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})}_{(b)},$ (F.3)
@@
Now, we are ready to prove the Lemma 1. By combining Lemma F.1, Lemma F.2, and Lemma F.3, we derive that
@@
where the second inequality holds because by setting $c=7\eta/6$ and $\lambda\geq\max\{84\sqrt{2}\eta,84d\eta\}$, we obtain:
@@
which conclude the proof of Lemma 1. ∎
Remark F.1.
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, then $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K\right)$.
F.2 Proofs of Lemmas for Lemma 1
F.2.1 Proof of Lemma F.1
Proof of Lemma F.1.
Let ~$\tilde{\ell}_{s}(\mathbf{w})=\ell_{s}(\mathbf{w}_{s})+\langle\mathbf{w}-\mathbf{w}_{s},\nabla\ell_{s}(\mathbf{w}_{s})\rangle+\frac{1}{2}\|\mathbf{w}-\mathbf{w}_{s}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s})}$ be a second-order approximation of the original function $\ell_{s}(\mathbf{w})$ at the point $\mathbf{w}_{s}$. The update rule (3) can also be expressed as
@@ F.2 Proofs of Lemmas for Lemma 1
Then, by Lemma F.4, we have
~$\langle\nabla\tilde{\ell}_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle\leq\frac{1}{2\eta}\left(\|\mathbf{w}_{s}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|_{H_{s}}^{2}-\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\|_{H_{s}}^{2}\right).$ (F.4)
To utilize Lemma F.6, we can rewrite the loss function as $\ell\left((x_{si}^{\top}\mathbf{w})_{i\in S_{s}},\mathbf{y}_{s}\right)=\ell_{s}(\mathbf{w})$. Consequently, according to Lemma F.6, it follows that
$\ell_{s}(\mathbf{w}_{s+1})-\ell_{s}(\mathbf{w}^{\star})\leq\langle\nabla\ell_{s}(\mathbf{w}_{s+1}),\mathbf{w}_{s+1}-\mathbf{w}^{\star}\rangle-\frac{1}{\zeta}\|\mathbf{w}_{s+1}-\mathbf{w}^{\star}\|^{2}_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})},$ (F.5)
where $\zeta=\log(K+1)+4$. Then, by combining (F.4) and (F.5), we have
@@
where in the second equality, we apply the Taylor expansion by introducing $\bm{\xi}_{s+1}$, a convex combination of $\mathbf{w}_{s+1}$ and $\mathbf{w}_{s}$. The first inequality follows from Lemma C.1 and Proposition C.1, the second inequality holds by Assumption 1, and the last inequality holds because
@@
where the third equality holds by setting $x_{s0}=\mathbf{0}$ for all $s\geq 1$.
@@
$\leq\zeta\left(\sum_{s=1}^{t}\ell_{s}(\mathbf{w}^{\star})-\sum_{s=1}^{t}\ell_{s}(\mathbf{w}_{s+1})\right)+4\lambda+6\sqrt{2}\zeta\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{2}^{2}-\sum_{s=1}^{t}\left\|\mathbf{w}_{s+1}-\mathbf{w}_{s}\right\|_{H_{s}}^{2},$
where the last inequality holds since $\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{H_{1}}^{2}\leq\lambda\|\mathbf{w}_{1}-\mathbf{w}^{\star}\|_{2}^{2}\leq 4\lambda$. Plugging in $\zeta=2\eta$, we conclude the proof. ∎
F.2.2 Proof of Lemma F.2
Proof of Lemma F.2.
@@
Note that $\tilde{\mathbf{z}}_{s}^{\mu}=\bm{\sigma}_{s}^{+}(\operatorname{smooth}^{\mu}_{s}(\bm{\sigma}_{s}(\tilde{\mathbf{z}}_{s})))$ by the definition of the pseudo inverse function $\bm{\sigma}_{s}^{+}$ such that $\bm{\sigma}_{s}^{+}(\bm{\sigma}_{s}(\mathbf{q}))=\mathbf{q}$ for any $\mathbf{q}\in\{\mathbf{p}\in[0,1]^{K_{s}}\mid\|\mathbf{p}\|_{1}<1\}$. Then, by Lemma F.7, we have
$\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s}^{\mu},\mathbf{y}_{s})-\sum_{s=1}^{t}\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq 2\mu t,\quad\text{and}\quad\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu).$ (F.6)
@@
$=\sum_{s=1}^{t}\langle\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s},\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\rangle-\sum_{s=1}^{t}\frac{1}{c_{\mu}}\|\mathbf{z}_{s}^{\star}-\tilde{\mathbf{z}}_{s}^{\mu}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2},$ (F.7)
where $c_{\mu}=\log(K+1)+2\log(1+(K+1)/\mu)+2$, the inequality holds by Lemma F.6, and the last equality holds by a direct calculation of the first order and Hessian of the logistic loss as follows:
@@
Then, one can easily verify that $\|\mathbf{d}^{\prime}_{s}\|_{\infty}\leq 1$ since $\|\mathbf{z}_{s}^{\star}\|_{\infty}\leq\max_{i\in S_{s}}\|x_{xi}\|_{2}\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|\tilde{\mathbf{z}}_{s}^{\mu}\|_{\infty}\leq\log(1+(K+1)/\mu)$. On the other hand, $\mathbf{d}_{s}^{\prime}$ is $\mathcal{F}_{s}$-measurable since $\mathbf{z}_{s}^{\star}$ and $\tilde{\mathbf{z}}_{s}^{\mu}$ are independent of $\mathbf{y}_{s}$. Moreover, we have $\|\mathbf{d}^{\prime}_{s}\|^{2}_{\mathbb{E}[\bm{\varepsilon}_{s}\bm{\varepsilon}_{s}^{\top}\mid\mathcal{F}_{s}]}=\|\mathbf{d}_{s}\|^{2}_{\mathbb{E}\left[(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})(\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s})^{\top}\mid\mathcal{F}_{s}\right]}=\|\mathbf{d}_{s}\|^{2}_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}$ and $\|\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})-\mathbf{y}_{s}\|_{1}\leq 2$. Thus, by Lemma F.5, with probability at least $1-\delta$, for any $t\geq 1$, we have
@@
$\leq(c_{\mu}+1)\sqrt{\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}}\cdot\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right),$ (F.8)
where the second inequality holds because $\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}=\mathbf{d}_{s}^{\top}\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})\mathbf{d}_{s}\leq 2$ and $\lambda\geq 1$. Then, combining (F.8) and (F.7), and rearranging the terms, we obtain